# When Text Embedding Meets Large Language Model: A Comprehensive Survey

### v3

@@ 4.2.1 Pooling Strategy
First pooling (only $\alpha_{1}=1$) is a common way of obtaining embedding for PLMs with bi-directional attention. For example, BERT splices the special token [CLS] before the input text, and the embedding output from the [CLS]’s position will be used as the embedding for the whole text [72]. Although T5-like models do not have the special token like [CLS] in BERT, Sentence-T5 [172] use the first embedding output by T5’s encoder and the [START] embedding output by T5’s decoder as the embedding of the whole text; however, these pooling strategies obtain slightly worse performance than mean pooling. For the decoder-only LLMs with causal attention, the first pooling does not work, as the embedding of the first position can’t be included in the semantics of the subsequent content.
@@
Last pooling (only $\alpha_{n}=1$) is a new pooling strategy emerging in the LLM era. Due to the uni-directionality of causal attention, only the last position shows information about the whole text. However, decoder-only LLMs are pre-trained by the next token prediction, and the embedding at the last position will align with the potential next token embedding [170]. Without additional interventions, there is no guarantee that the embedding contains the semantics of the entire text. Three common interventions here are
  • Prompt-Based Last Pooling. A special prompt is introduced to induce the model to summarize the semantics of the whole text in the last position [170, 176, 178, 184, 191, 192]. For example, PromptEOL [170] propose a prompt template The sentence [X] means in a word:‘‘, where [X] is a placeholder. In practice, [X] is filled with the input text, and the last pooling strategy is used to obtain the text embedding. Although there are many variants, the core idea in the line of work is consistent, i.e., convert text embedding to language modeling and guide the model to summarize the whole text semantics in the last position. These methods are marked as “last” in the “pooling” column of Table 4.
  • Special-Token Last Pooling. Some works [195, 196, 180, 184, 183, 182, 181] introduce a special token, such as <EOS>, to obtain embedding in the last position, while the LLM is fine-tuned incrementally, learning to converge semantics at the position of the special token. These methods are marked as “special last” in the “pooling” column of Table 4.
  • Special-Sequence Last Pooling. ChatRetriever [183] and DEBATER [188] add a special token sequence, such as $[\texttt{EMB}_{1}],\cdots,[\texttt{EMB}_{t}]$, at the end of the input text. The authors argue that these $t$ consecutive special tokens serve as a chain of thought that extends and guides the learning space for more effective embeddings. These methods are marked as “special seq last” in the “pooling” column of Table 4.
Partial Pooling
Partial Pooling ($\alpha_{i}=1$ for $i$ is greater than a specific $k$) is also an effective method to mitigate the effects of causal attention, which requires prompt-based assistance. Echo [190] propose a prompt: Rewrite the sentence:[x], rewritten sentence:[x], where [x] is the placeholder. In practice, both placeholders are filled with the same text, and the mean pooling strategy is used to obtain the text embedding, but it is pooled only within the range of the second occurrence of the text. In this way, the entire text is already present in the first placeholder, so each token populated in the second placeholder can access the entire text information through causal attention. Mistral-SPLADE [186] follows Echo’s approach and uses the obtained embedding to generate sparse representations.
@@
Some works add complex modules with attention mechanisms before the conventional pooling strategy. NV-Embed [199] introduces an additional attention-based network before mean pooling to post-interact with the hidden states $\mathcal{H}$. The attention-based network consists of a cross-attention block and an MLP, where the key and value matrix used in the cross-attention are learnable and low-rank, and $\mathcal{H}$ is considered as the query. Mistral-SPLADE [186] uses the same prompt to Echo [190] to obtain text embeddings and follow the SPLADE recipe [205] to fine-tune the Mistral-based embedded. LENS [204] further clusters the LLM’s tokens according to their embedding to narrow the sparse representation latitude and reduce information redundancy, achieving better sparsification results.
@@
Some works use post-interactive modules to mix the information from the different layers’ hidden states in the LLM. For example, [200] proposes a trainable multi-layer pooling method that aggregates text embeddings from multiple layers using last/mean pooling. These embeddings form a matrix $\mathcal{H}^{L}$, which is processed by a cross-attention Transformer. Cross-attention is computed by summing $\mathcal{H}^{L}$ with a learnable weight matrix, then deriving key and value representations. A fixed learnable query is used across inputs, and the feed-forward block output serves as the final embedding. [185] collect the hidden states from the layer with the optimal alignment and uniformity metric [206], which denotes $\mathcal{H}^{a}$ and $\mathcal{H}^{u}$ separately. Then, a multi-layer network called LMORT is introduced to fuse the information in both $\mathcal{H}^{a}$ and $\mathcal{H}^{u}$, where each layer consists of a bi-directional self-attention block, a bi-directional cross-attention block, a feed-forward block, and inter-block residual mechanism. The authors show that the performance of retrieval tasks can be improved by only fine-tuning the LMORT’s parameters. MoEE [192] use the routing weights from Mixture-of-Experts (MOE) LLMs as the embeddings, showing a promising complementarity of hidden states from last layers. ReBA [193] aggregates the weights of all attention heads in different layers to recompute the final token embedding and mitigates the shortcomings of causal attention by utilizing a similar “repetition” idea for Echo [193].
@@ 4.2.2 Attention Mechanism
The discussion of attention in the current approach focuses on decoder-only LLM-based embedder. Causal Attention [207] is a common operation for decoder-based LLMs, which ensures that the language modeling can only refer to the prefix to predict the next token. However, extensive empirical studies [208, 178, 190] have shown that causal attention will degrade downstream task performance. Current approaches either (1) retain the causal attention and use other tricks (e.g., special pooling methods) or (2) convert it to bi-directional attention and make the model adapt to it.
Maintain Causal Attention
If the model maintains causal attention, it is often necessary to resort to pooling tricks to mitigate the resulting negative effects. Representatives of these methods include SGPT [173], PromptEOL [170], and Echo [190]; please refer to Section 4.2.1 for details.
Convert to Bi-Directional Attention
Considering that the causal attention is implemented by a mask matrix implementation in Transformer  [^10] Transformers (https://github.com/huggingface/transformers) is the most popular open-source toolkit to instantiate LLMs., it is convenient and quick to remove the mask matrix during incremental fine-tuning. BeLLM [178] pioneered an attempt to change the causal attention of the last few layers in the decoder-only LLM to bi-directional attention. This trial is motivated by the fact that as the number of layers increases, the performance of downstream tasks does not improve monotonically, but there is a turning point. When fully converting the causal attention in decoder-only LLM to bi-directional, sufficient data or additional training tasks are required to adapt the LLM to the new attentional mechanism. The experimental results in [178] show that when training on small-scale data, changing the attention of only the last layer from causal to bi-directional can effectively improve the performance on the SemEval Benchmark; however, changing all the attentional layers directly to bi-directional will result in a significant decrease in performance. Further, NV-Embed [199] shows that LLMs can adaptively update the parameters and complete the conversion from causal attention to bi-attention without additional operations when using sufficient data (hundreds of datasets) for incremental fine-tuning. To adapt the LLM to use bi-direction attention without massive training data, LLM2Vec [168] propose a self-supervised task, masked next token prediction (MNTP). MNTP combines the ideas of mask language modeling and next token prediction, where the token is first randomly masked off from the text, and the hidden state from the previous position of the mask token is used to predict the mask token.
@@ 4.2.3 Additional Projector
Adding projectors is the usual strategy in text embedding. Unlike pooling strategies that operate on multiple hidden states, projectors usually operate on a single embedding for other purposes. We divide the existing work into three parts according to the purpose of the projectors.
@@
The hidden layer dimension (width) usually increases with the number of total parameters, according to the scaling laws [209]. The hidden states of T5-11B have 1024 dimensions, while the number is 4096 for the mainstream 7B decoder-only LLMs, which are both higher than the 768 dimensions for a bunch of BERT-based models. In practice, high-dimensional embedding dramatically increases the resource overhead for storage and inference, and a simple approach is to add a projector $g:\mathbb{R}^{d}\to\mathbb{R}^{m},m<d$ after pooling to obtain the low-dimensional embedding. With a fixed output dimension linear layer, Sentence-T5 [172] and GTR [179] empirically demonstrate that the performance on STS, classification, and retrieval tasks can be improved by scaling the embedder’s parameter only without scaling dimension. GirtLM [169] also uses a linear layer to map Mistral’s 4096-dimensional embedding to 1024 and finds that the performance of MTEB would be slightly degraded.
@@
Similar to the purpose of dimensionality reduction, converting text embedding to sparse representations can improve the inference efficiency of the model [205] and enhance its performance on partial downstream tasks such as long document retrieval [56]. For BERT-based embedders, text embeddings can be mapped to the vocabulary-length logits through the projector for mask language modeling; then, the vocabulary-size logits can be regarded as the text representation. To sparsify the representation, common methods include gate mechanism [210], top-k masking [211], and regularization terms [212]. Similarly, text embeddings from LLM-based embedders can go through the decoder layer and use a similar idea for sparsification. Specifically, PromptReps [184] uses the prompt similar to PromptEOL to obtain text embeddings and introduce ReLU function, log-saturation [205] and top-k masking to obtain the sparse representation for documents. In addition, an empirical study [213] finds that the embedding output from LLM-based embedders, which is fine-tuned by contrastive learning, still produces high-quality sparse representations via the decoding layer and simple top-$k$ masking.
Projector for Other Purposes
Search-Adaptor [214] proposes the use of additional projectors for customizing commercial text embedding interfaces, improving performance on private domain data.
@@ 4.3.1 Training Free (TF)
Some works leverage the inherent capabilities of large language models without additional fine-tuning.
Prompt Engineering
@@
By carefully crafting input prompts, practitioners can guide the model to produce desirable embeddings tailored to specific tasks. In the era of PLM, PromptBERT [217] first proposed a prompt-based method to enhance the text representation of the BERT model in training-free setting. Specifically, this method uses a fixed template (e.g., "[X] is [mask]", where [X] is a placeholder.) and leverages the hidden states at the mask position as the sentence representation. Recently, two lines of methods for LLMs have been developed:
@@
  • Summary-Style Prompt. Several approaches leverage prompt engineering to condense input text into a single word, aiming to capture core semantics. For instance, PromptEOL [170] employs the template ‘‘The sentence [X] means in a word:’’ to elicit a one-word summary, where [X] is replaced with the input text. The final hidden state is then used as the sentence embedding. Furthermore, PromptEOL enhances performance by utilizing in-context learning, generating one-word summaries with GPT-4 for a sentence and prepending them to the input as contextual examples. Building on this, PromptReps [184] expands the utility of prompt engineering by not only crafting dense text embeddings but also enabling sparse bag-of-words representations derived from language models. To further enhance LLM embedding expressiveness, PromptSTH/SUM [176] proposes a Pretended Chain of Thought and Knowledge Enhancement. Pretended Chain of Thought simulates step-by-step reasoning by prepending prompts with phrases like "After thinking step by step," encouraging detailed sentence representation. Knowledge Enhancement integrates human summarization expertise by incorporating relevant knowledge into prompts, guiding the model to focus on essential information. MetaEOL [175] guides LLMs to produce embeddings through a series of carefully designed prompts that address multiple representational aspects, then use an average of these meta-task-derived embeddings as the final representation. Similarly, GenEOL [191] introduces a novel approach by generating diverse sentence variations of the target text using an LLM. Subsequently, it computes the final sentence embedding by averaging the embeddings of these generated variants. Different from these, MoEE [192] leverages the inherent structure of Mixture-of-Experts (MoE) LLMs, combining routing weights and hidden states to produce robust embeddings without requiring explicit fine-tuning.
@@
  • Repetition-Style Prompt. The inherent causal attention mechanism in decoder-only Large Language Models (LLMs) can restrict information flow for embedding models, prompting research into mitigation strategies via prompt design. Echo et al. [190] introduce a self-repetition prompt, Rewrite the sentence:[x], rewritten sentence:[x], where [x] represents the input sequence. This method leverages the repetition of the input to enhance contextual understanding, extracting text embeddings through mean pooling applied solely to the second occurrence of [x]. ReBA [193] extends this concept by developing an algorithm that directly modifies the model’s attention matrix, using the repeated input to refine hidden state representations at each position, leading to enhanced performance compared to basic repetition. Token Prepending [177] proposes a plug-and-play, training-free approach. This technique involves prepending the decoded sentence embedding from each layer to the input of the subsequent layer. By doing so, it allows earlier tokens to access the complete sentence embedding, effectively circumventing the limitations imposed by causal attention.

### v4

@@ 4.2.1 Traditional Pooling Strategy
First pooling is a common way of obtaining an embedding for PLMs with bi-directional attention, which can be express as
${\bm{h}}=\operatorname{\mathcal{P}}([{\bm{h}}^{(L)}_{0},{\bm{h}}^{(L)}_{x_{1}},\cdots{\bm{h}}^{(L)}_{x_{n}},{\bm{h}}^{(L)}_{s_{1}},\cdots,{\bm{h}}^{(L)}_{s_{k}}])={\bm{h}}^{(L)}_{0}$ (21)
For example, BERT splices the special token [CLS] before the input text, and the embedding output from the [CLS]’s position will be used as the embedding for the whole text [72]. Although T5-like models do not have the special token like [CLS] in BERT, Sentence-T5 [168] use the first embedding output by T5’s encoder and the [START] embedding output by T5’s decoder as the embedding of the whole text; however, these pooling strategies obtain slightly worse performance than mean pooling. For the decoder-only LLMs with causal attention, the first pooling does not work, as the embedding of the first position can’t be included in the semantics of the subsequent content.
Mean Pooling
Traditional Mean pooling averages the embedding of each position and shows better performance in BERT than first pooling [72, 168].
${\bm{h}}=\operatorname{\mathcal{P}}([{\bm{h}}^{(L)}_{x_{1}},\cdots{\bm{h}}^{(L)}_{x_{n}},{\bm{h}}^{(L)}_{s_{1}},\cdots,{\bm{h}}^{(L)}_{s_{k}}])=\sum_{i=1}^{n}\alpha_{i}{\bm{h}}_{x_{i}}+\sum_{j=1}^{k}\beta_{j}{\bm{h}}_{s_{j}}$ (22)
  • Naive Mean Pooling (Mean) The naive mean pooling generally does not include special tokens and $\forall i,j$ satisfies $\alpha_{i}=1/n,\ \beta_{j}=0$ in Eqn 20. Up to now,, mean pooling remains regarded as one of the most straightforward methods, widely employed across various LLM backbones.
  • Weighted Mean Pooling (Weighted Mean) The weighted mean pooling fits $\alpha_{i}=i/\sum_{i=1}^{n}i,\ \beta_{j}=0$ for $\forall i,j$ in Eqn 22. For decoder-based LLMs, assigning the same weight to each position because the latter position will be able to see more semantic information due to causal attention [176]. SGPT [176] suggests the use of weighted mean pooling with the intuition that the latter position should be given a larger weight. L3Prune [200] follows SGPT’s pooling strategy to fine-tune the pruned LLM-based embedders.
  • Partial Mean Pooling (Partial Mean) The partial mean pooling assign $\alpha_{i}=1/l$ for $\forall i\in[k,k+l]$ while the other elements is 0 in Eqn 22. $k$ is an example-specific number and $l$ is the token number of the input text $x$. This strategy is viewed as an effective method to mitigate the effects of causal attention, which requires prompt-based assistance. Echo [189] propose a prompt: Rewrite the sentence:[x], rewritten sentence:[x], where [x] is the placeholder. In practice, both placeholders are filled with the same text, and the mean pooling strategy is used to obtain the text embedding, but it is pooled only within the range of the second occurrence of the text. In this way, the entire text is already present in the first placeholder, so each token populated in the second placeholder can access the entire text information through causal attention. Mistral-SPLADE [183] follows Echo’s approach and uses the obtained embedding to generate sparse representations.
  • Special-Sequence Mean Pooling (Special Seq Mean) The special-sequence mean pooling can be denoted as $\alpha_{i}=0,\beta_{j}=1/k$ for $\forall i,j$. in Eqn.22. AutoRegEmbed [174] is the only method employing this pooling strategy, which draws inspiration from long text compression techniques. The authors introduce the document prediction task, guiding the embedder condenses the semantic information of the entire text into multiple special tokens. Then the mean pooling on all special tokens aggregates the information from these tokens to obtain the text embedding.
@@
Last pooling is a new pooling strategy emerging in the LLM era. Due to the unidirectionality of causal attention, only the last position shows information about the whole text. However, decoder-only LLMs are pre-trained by the next token prediction, and the embedding at the last position will align with the potential next token embedding [169]. Without additional interventions, there is no guarantee that the embedding contains the semantics of the entire text. Therefore, last pooling ultimately requires the use of prompts and/or special tokens to achieve semantic aggregation, which can be expressed as follows in the technique:
[…]
  • Prompt-Based Last Pooling (Last). A special prompt is introduced to induce the model to summarize the semantics of the whole text in the last position [169, 171, 173, 181, 190, 191]. For example, PromptEOL [169] propose a prompt template The sentence [X] means in a word:‘‘, where [X] is a placeholder. In practice, [X] is filled with the input text, and the last pooling strategy is used to obtain the text embedding. Although there are many variants, the core idea in the line of work is consistent, i.e., convert text embedding to language modelling and guide the model to summarize the whole text semantics in the last position. These methods are marked as “last” in the “pooling” column of Table 4.
  • Special-Token Last Pooling (Special Last). Some works [194, 195, 177, 181, 180, 179, 178] introduce a special token, such as <EOS>, to obtain embedding in the last position, while the LLM is fine-tuned incrementally, learning to converge semantics at the position of the special token.
  • Special-Sequence Last Pooling (Special Seq Last). ChatRetriever [180], DEBATER [185] and AutoRegEmbed [174] add a special token sequence, such as $\texttt{EMB}_{1}],\cdots,[\texttt{EMB}_{t}]$, at the end of the input text. The authors argue that these $t$ consecutive special tokens serve as a chain of thought that extends and guides the learning space for more effective embeddings.
@@
Some works add complex modules with attention mechanisms before the conventional pooling strategy, where $\mathcal{P}(\cdot)$ is a learnable network. NV-Embed [217] introduces an additional attention-based network before mean pooling to post-interact with the hidden states $\mathcal{H}$. The attention-based network comprises a cross-attention block and an MLP, where the key and value matrices used in the cross-attention are learnable and low-rank, and $\mathcal{H}$ is considered the query. Mistral-SPLADE [183] uses the same prompt as Echo [189] to obtain text embeddings and follows the SPLADE recipe [224] to fine-tune the Mistral-based embedding. LENS [201] further clusters the LLM’s tokens according to their embedding to narrow the sparse representation latitude and reduce information redundancy, achieving better sparsification results.
4.2.2 Special Pooling Strategy
Some novel approaches explore methods that do not (solely) rely on the hidden states of the final layer. Therefore, they cannot be expressed using Eqn 20. We categorize these methods into two types: (1) Multi-layer Pooling, which attempts to enhance embeddings by utilizing information contained in outputs from other layers of the model; (2) Generative Pooling, which seeks to incorporate the content generated by LLMs into the pooling process. Below, we will provide a detailed introduction to each method:
@@ 4.2.2 Special Pooling Strategy
Some works mix the information from the different layers’ hidden states in the LLM. For example, MLTP [198] proposes a trainable multi-layer pooling method that aggregates text embeddings from multiple layers using last/mean pooling. These embeddings form a matrix $\mathcal{H}^{L}$, which is processed by a cross-attention Transformer. Cross-attention is computed by summing $\mathcal{H}^{L}$ with a learnable weight matrix, then deriving key and value representations. A fixed learnable query is used across inputs, and the feed-forward block output serves as the final embedding. [182] collects the hidden states from the layer with the optimal alignment and uniformity metric [225], which denotes $\mathcal{H}^{a}$ and $\mathcal{H}^{u}$ separately. Then, a multi-layer network called LMORT is introduced to fuse the information in both $\mathcal{H}^{a}$ and $\mathcal{H}^{u}$, where each layer consists of a bi-directional self-attention block, a bi-directional cross-attention block, a feed-forward block, and an inter-block residual mechanism. The authors demonstrate that the performance of retrieval tasks can be enhanced by fine-tuning the LMORT’s parameters alone. MoEE [191] utilizes the routing weights from Mixture-of-Experts (MOE) LLMs as embeddings, demonstrating a promising complementarity with hidden states from the last layers. ReBA [192] aggregates the weights of all attention heads in different layers to recompute the final token embedding while utilizing a similar “repetition” idea proposed by Echo [192].
Generative Pooling
Some works enhance embedding quality through multi-step reasoning or augment the original input using generated components; therefore, they incorporate the generated part by LLMs into pooling strategies.
  • Generated Sequence Mean Pooling (Gen. Seq Mean) GIRCSE [206] sets an iterative step count $k$, and uses the average of the final hidden states at the generated $k$ positions as the text embedding.
  • Generated Sequence Last Pooling (Gen. Seq Last) O1 Embedder [186] and Search-R3 [187] employ supervised fine-tuning and reinforcement learning approaches respectively to enable the LLM to perform reasoning first, ultimately generating a special token to obtain the embedding.
  • Hybrid Sequence Mean (Hyb. Seq Mean) GRACE [206] averages the last hidden states of both input and output content, excluding the instruction text, which is regarded as the text embedding.
@@ 4.2.3 Attention Mechanism
The discussion of attention in the current approach focuses on a decoder-only LLM-based embedder. Causal Attention [226] is a common operation for decoder-based LLMs, which ensures that the language modeling can only refer to the prefix to predict the next token. However, extensive empirical studies [227, 173, 189] have shown that causal attention will degrade downstream task performance. Current approaches either (1) retain the causal attention and use other tricks (e.g., special pooling methods) or (2) convert it to bi-directional attention and make the model adapt to it.
Causal Attention
If the model maintains causal attention, it is often necessary to resort to pooling tricks to mitigate the resulting adverse effects. Representatives of these methods include SGPT [176], PromptEOL [169], and Echo [189]; please refer to Section 4.2.1 for details.
Bi-Directional Attention
Considering that the causal attention is implemented by a mask matrix implementation in Transformer  [^10] Transformers (https://github.com/huggingface/transformers) is the most popular open-source toolkit to instantiate LLMs., it is convenient and quick to remove the mask matrix during incremental fine-tuning. BeLLM [173] pioneered an attempt to change the causal attention of the last few layers in the decoder-only LLM to bi-directional attention. This trial is motivated by the fact that, as the number of layers increases, the performance of downstream tasks does not improve monotonically, but there is a turning point. When fully converting the causal attention in decoder-only LLM to bi-directional, sufficient data or additional training tasks are required to adapt the LLM to the new attentional mechanism. The experimental results in [173] show that when training on small-scale data, changing the attention of only the last layer from causal to bi-directional can effectively improve performance on the SemEval Benchmark; however, directly changing all the attentional layers to bi-directional results in a significant decrease in performance. Furthermore, NV-Embed [217] demonstrates that LLMs can adaptively update their parameters and complete the conversion from causal attention to bi-attention without requiring additional operations, provided sufficient data (hundreds of datasets) is available for incremental fine-tuning. To adapt the LLM to use bi-directional attention without massive training data, LLM2Vec [193] proposes a self-supervised task, masked next token prediction (MNTP). MNTP combines the ideas of mask language modeling and next token prediction, where the token is first randomly masked from the text, and the hidden state from the previous position of the mask token is used to predict the masked token. In addition, Conan-Embedding-v2 [210] proposes a novel approach: smoothly transitioning from causal attention to bidirectional attention during training.
@@ 4.2.4 Additional Projector
Adding projectors is the usual strategy in text embedding. Unlike pooling strategies that operate on multiple hidden states, projectors usually operate on a single embedding for other purposes. We divide the existing work into three parts based on the projectors’ purposes.
@@
The hidden layer dimension (width) usually increases with the number of total parameters, according to the scaling laws [228]. The hidden states of T5-11B have 1024 dimensions, whereas the mainstream 7B decoder-only LLMs have 4096 dimensions, both of which are higher than the 768 dimensions of a number of BERT-based models. In practice, high-dimensional embedding dramatically increases the resource overhead for storage and inference, and a simple approach is to add a projector $g:\mathbb{R}^{d}\to\mathbb{R}^{m},m<d$ after pooling to obtain the low-dimensional embedding. With a fixed output dimension linear layer, Sentence-T5 [168] and GTR [175] empirically demonstrate that the performance on STS, classification, and retrieval tasks can be improved by scaling the embedder’s parameter only without scaling the dimension. GirtLM [197] also uses a linear layer to map Mistral’s 4096-dimensional embedding to 1024 and finds that the performance of MTEB would be slightly degraded.
@@
Similar to the purpose of dimensionality reduction, converting text embedding to sparse representations can improve the inference efficiency of the model [224] and enhance its performance on partial downstream tasks such as long document retrieval [56]. For BERT-based embedders, text embeddings can be mapped to the vocabulary-length logits through the projector for mask language modeling; then, the vocabulary-size logits can be regarded as the text representation. To sparsify the representation, common methods include gate mechanisms [229], top-k masking [230], and regularization terms [231]. Similarly, text embeddings from LLM-based embedders can go through the decoder layer and use a similar idea for sparsification. Specifically, PromptReps [181] uses the prompt similar to PromptEOL to obtain text embeddings and introduce ReLU function, log-saturation [224] and top-k masking to obtain the sparse representation for documents. In addition, an empirical study [232] finds that the embedding output from LLM-based embedders, which is fine-tuned by contrastive learning, still produces high-quality sparse representations via the decoding layer and simple top-$k$ masking.
Projector for Customized Adaptation
Search-Adaptor [233] proposes the use of additional projectors for customizing commercial text embedding interfaces, improving performance on private domain data. Furthermore, Matryoshka-Adaptor [234] and SMEC [235] attempt to obtain embeddings across different dimensions while customizing the adaptation, and maintain the quality of low-dimensional embeddings by employing distinct optimization objectives.
@@ 4.3.1 Training Free (TF)
Some works leverage the inherent capabilities of large language models without additional fine-tuning. By carefully crafting input prompts, practitioners can guide the model to produce embeddings that are desirable and tailored to specific tasks. In the era of PLM, PromptBERT [238] first proposed a prompt-based method to enhance the text representation of the BERT model in training-free setting. Specifically, this method uses a fixed template (e.g., "[X] is [mask]", where [X] is a placeholder) and leverages the hidden states at the mask position as the sentence representation. Recently, two lines of methods for LLMs have been developed:
@@
  • Summary-Style Prompt. Several approaches leverage prompt engineering to condense input text into a single word, aiming to capture core semantics. For instance, PromptEOL [169] employs the template ‘‘The sentence [X] means in a word:’’ to elicit a one-word summary, where [X] is replaced with the input text. The final hidden state is then used as the sentence embedding. Furthermore, PromptEOL enhances performance by utilizing in-context learning, generating one-word summaries with GPT-4 for a sentence and prepending them to the input as contextual examples. Building on this, PromptReps [181] expands the utility of prompt engineering by not only crafting dense text embeddings but also enabling sparse bag-of-words representations derived from language models. To further enhance LLM embedding expressiveness, PromptSTH/SUM [171] proposes a Pretended Chain of Thought and Knowledge Enhancement. Pretended Chain of Thought simulates step-by-step reasoning by prepending prompts with phrases like "After thinking step by step," encouraging detailed sentence representation. Knowledge Enhancement integrates human summarization expertise by incorporating relevant knowledge into prompts, guiding the model to focus on essential information. MetaEOL [170] guides LLMs to produce embeddings through a series of carefully designed prompts that address multiple representational aspects, then uses an average of these meta-task-derived embeddings as the final representation. Similarly, GenEOL [190] introduces a novel approach by generating diverse sentence variations of the target text using an LLM. Subsequently, it computes the final sentence embedding by averaging the embeddings of these generated variants. Different from these, MoEE [191] leverages the inherent structure of Mixture-of-Experts (MoE) LLMs, combining routing weights and hidden states to produce robust embeddings without requiring explicit fine-tuning.
@@
  • Repetition-Style Prompt. The inherent causal attention mechanism in decoder-only LLMs can restrict information flow for embedding models, prompting research into mitigation strategies via prompt design. Echo et al. [189] introduce a self-repetition prompt, Rewrite the sentence:[x], rewritten sentence:[x], where [x] represents the input sequence. This method leverages the repetition of the input to enhance contextual understanding, extracting text embeddings through mean pooling applied solely to the second occurrence of [x]. ReBA [192] extends this concept by developing an algorithm that directly modifies the model’s attention matrix, using the repeated input to refine hidden state representations at each position, leading to enhanced performance compared to basic repetition. Token Prepending [172] proposes a plug-and-play, training-free approach. This technique involves prepending the decoded sentence embedding from each layer to the input of the subsequent layer. By doing so, it allows earlier tokens to access the complete sentence embedding, effectively circumventing the limitations imposed by causal attention.